# Monero’s Decentralized P2P Exchanges: Functionality, Adoption, and Privacy Risks

### v2

@@ 3.3 BasicSwap
Network Architecture: BasicSwap leverages the SMSG network as its underlying P2P network. SMSG functions as a decentralized storage network where each node maintains a temporary copy of every end-to-end encrypted message broadcast within the last 48 hours. Every BSX client operates a SMSG node serving the broader Particl P2P infrastructure that supports all Particl decentralized applications. A front-end is hosted locally with every running BSX client. Within this interface, users can interact with posted trade offers. The trade offers are broadcast using the SMSG network, enabling each peer to maintain its own copy of the decentralized order book without relying on centralized matching services. The DEX trading functionality is implemented through atomic swap protocols, with the SMSG network handling the necessary communication required to execute these trustless swaps. Trade Protocol: BasicSwap deploys two atomic swap protocols. If both coins support Bitcoin-style scripting, Hash Time-Locked Contracts (HTLCs) [7] are used, enforcing swaps through on-chain scripts with shared secret hashes and timelocks. If a coin like Monero lacks these capabilities, swaps rely on adaptor signatures [11], where private key shares and one-time verifiably encrypted signatures ensure atomicity. Users create and accept offers via BSX’s graphical interface, while the application automates the swap process, handling secrets, keys, and transaction ID sharing via the SMSG network. Trade Fee: No additional trade fees apply, only network transaction fees. Dispute Resolution: Both protocols ensure swaps either complete or refund both parties without third-party intervention.
Governance: BasicSwap DEX is built within the Particl network, governed as a decentralized autonomous organization. Details on how the DAO governs BSX is not provided but as trades do not include any fee related to the DAO or another third-party, financial incentives and involvement in the project appear to be based on community participation rather than direct revenue generation.
Operational Details: Running blockchain nodes for multiple cryptocurrencies is needed and creates substantial resource requirement. While Bitcoin synchronization can be accelerated using snapshots and Monero can utilize remote RPC endpoints, the initial setup remains time-consuming. Concurrent node operation demands significant computation, storage, and network resources, potentially deterring casual and test users. These known constraints are scheduled to be addressed as the DEX progresses beyond its current beta phase.

### v3

@@ 3.3 BasicSwap
Network Architecture: BasicSwap leverages the SMSG network as its underlying P2P network. SMSG functions as a decentralized storage network where each node maintains a temporary copy of every end-to-end encrypted message broadcast within the last 48 hours. Every BSX client operates a SMSG node serving the broader Particl P2P infrastructure that supports all Particl decentralized applications. A front-end is hosted locally with every running BSX client. Within this interface, users can interact with posted trade offers. The trade offers are broadcast using the SMSG network, enabling each peer to maintain its own copy of the decentralized order book without relying on centralized matching services. The DEX trading functionality is implemented through atomic swap protocols, with the SMSG network handling the necessary communication required to execute these trustless swaps. Trade Protocol: BasicSwap deploys two atomic swap protocols. If both coins support Bitcoin-style scripting, Hash Time-Locked Contracts (HTLCs) [7] can be used, enforcing swaps through on-chain scripts with shared secret hashes and timelocks. If a coin like Monero lacks these capabilities, swaps rely on adaptor signatures [11], where private key shares and one-time verifiably encrypted signatures ensure atomicity. Users create and accept offers via BSX’s graphical interface, choosing which coins to trade and which swap protocol to use when the involved coins support both. The application then automates the swap process, handling secrets, keys, and transaction ID sharing via the SMSG network. Trade Fee: No additional trade fees apply, only network transaction fees. Dispute Resolution: Both protocols ensure swaps either complete or refund both parties without third-party intervention.
Governance: The Particl network applies a DAO structure [21, DAO and Network Treasury], details on how the DAO governs BSX is not provided but as trades do not include any fee related to the DAO or another third-party. Financial incentives and involvement in the project appear to be independent from Particl based on community participation rather than direct revenue generation.
Operational Details: Running blockchain nodes for multiple cryptocurrencies is needed and creates substantial resource requirement. Although Bitcoin synchronization can be accelerated using snapshots and Monero can utilize remote RPC endpoints, the initial setup remains time-consuming. Concurrent node operation demands significant computation, storage, and network resources, potentially deterring casual and test users. These known constraints are scheduled to be addressed as the DEX progresses beyond its current beta phase.